OpenStudy (anonymous):

Assume f (x) = -2x+8 and g (x) =3x. what is the value of (f+g) (-1)?

OpenStudy (isaiah.feynman):

First what is (f+g)(x)?

OpenStudy (anonymous):

f (x) + g (x)?

OpenStudy (kc_kennylau):

(f+g)(-1)=f(-1)+g(-1)=plz continue :)

OpenStudy (kc_kennylau):

yep :D

OpenStudy (isaiah.feynman):

Yes. That means add those two expressions.

OpenStudy (shamil98):

\[f(x) = -2x + 8\] Evaluate f(-1) and \[g(x) = 3x\] at g(-1) then add those two results together.

OpenStudy (zzr0ck3r):

another way f (x) = -2x+8 and g (x) =3x. (f+g)(x) = -2x+8 + 3x = x+8 so (f+g)(-1) = -1+8

OpenStudy (anonymous):

so in the end it'd be 7? :)

OpenStudy (shamil98):

yup

OpenStudy (zzr0ck3r):

did you just ask if -1+8 = 7?

OpenStudy (zzr0ck3r):

:P

OpenStudy (isaiah.feynman):

@zzr0ck3r that's my way. :P